# Tractable Algorithms for Changepoint Detection in Player Performance Metrics

### v1

@@ Abstract
We present tractable methods for detecting changes in player performance metrics and apply these methods to Major League Baseball (MLB) batting and pitching data from the 2023 and 2024 seasons. First, we derive principled benchmarks for when performance metrics can be considered statistically reliable, assuming no underlying change, using distributional assumptions and standard concentration inequalities. We then propose a changepoint detection algorithm that combines a likelihood-based approach with split-sample inference to control false positives, using either nonparametric tests or tests appropriate to the underlying data distribution. These tests incorporate a shift parameter, allowing users to specify the minimum magnitude of change to detect. We demonstrate the utility of this approach across several baseball applications: detecting changes in batter plate discipline metrics (e.g., chase and whiff rate), identifying velocity changes in pitcher fastballs, and validating velocity changepoints against a curated ground-truth dataset of pitchers who transitioned from relief to starting roles. Our method flags meaningful changes in 91% of these ‘ground-truth’ cases and reveals that, for some metrics, more than 60% of detected changes occur in-season. While developed for baseball, the proposed framework is broadly applicable to any setting involving monitoring of individual performance over time.
@@ 1 Introduction
Aaron Judge, an outfielder for the New York Yankees, led Major League Baseball (MLB) in on-base percentage (OBP) during the 2024 season, finishing with an impressive OBP of .458 compared to the league average of .312. Early in the season, however, his OBP after his first 24 plate appearances (PA) was just .208, substantially lower than his end-of-season OPB. Intuitively, this discrepancy highlights the well-known issue that performance metrics based on small sample sizes can vary substantially, and that this variability typically decreases as sample size increases. This example naturally raises a fundamental question relevant not just in sports analytics but in any field that tracks performance over time: how can we efficiently detect meaningful changes in a metric, or determine when apparent changes are simply due to small sample sizes?
@@
Across sports, sample size in relation to metric stabilization and changepoints is a widely discussed topic with researchers often asking how large a sample is needed to be reasonably confident in the corresponding results (Carleton 2012, Foley 2024). This concern about small sample variability is foundational in statistics, with the James-Stein estimator providing a classic example in baseball: even seemingly straightforward (batting) averages can be misleadingly variable when based on limited data (Efron & Morris 1977). In this paper, we formalize the concept of metric stabilization, proposing methods for determining when a metric becomes sufficiently stable in baseball, and introduce algorithms for detecting changes in player performance metrics over time.
We note that calculations to determine metric stabilization points should not be entirely based on player data (especially spanning multiple seasons), as players often change over the course of the season, as well as across seasons. In fact, our changepoint analysis (see Sections 4-6) indicates that for some statistics more than 60% of detected changes occur outside of the first or last month of the season (i.e., occuring in May through August).
@@
We present efficient algorithms for detecting changepoints in both binary and continuous performance metrics at the individual player level. While changepoint detection methods have been used in prior sports analytics research, they have primarily focused on identifying changes in aggregated statistics for the average player, team, or league across seasons (Chiou et al. 2016, Whalen et al. 2024). In contrast, our focus is on identifying changes in individual performance across time, which poses a unique computational challenge due to the large number of players to monitor in real time.
Our approach draws on ideas from the changepoint detection literature (Hinkley & Hinkley 1970, Sen & Srivastava 1975, Killick & Eckley 2014, Matteson & James 2014, Shang et al. 2017, Jewell et al. 2022) but integrates them in a novel framework specifically designed for this application. In particular, we incorporate split-sample inference with hypothesis tests that include a shift parameter, giving users direct control over both the Type I error rate and the magnitude of changes considered meaningful. This addresses a key limitation in existing R packages (Lindeløv 2024), which typically assume normally distributed data and omit split-sample inference, and therefore tend to overflag changepoints and perform poorly for binary outcomes like plate discipline metrics. We provide open-source R functions that handle both continuous and binary data efficiently, using a likelihood ratio statistic maximized across all possible changepoints for a given player. We validate our method on a constructed ground-truth dataset of pitchers who changed roles from reliever to starter during the 2023 or 2024 MLB seasons, and illustrate its broader utility by detecting changes in fastball velocity and plate discipline metrics such as zone swing percentage, whiff percentage, and chase rate.
Although this paper focuses on applications to MLB, the algorithms and software we present are broadly applicable to any domain where detecting changes in individual performance metrics is of interest, such as in manufacturing, where changepoint detection has been a longstanding area of research (Shang et al. 2017).
@@
The paper is organized as follows. Section 2 describes the MLB datasets. In Section 3, we formalize methods for benchmarking metric stability under the assumption of no changepoints and present stabilization points for several common MLB batting statistics. Section 4 introduces our changepoint detection algorithms. We apply these algorithms to batter plate discipline metrics in Section 5, and to pitcher fastball velocity in Section 6. Ground-truth validation is presented in Section 7. Section 8 concludes with a discussion of limitations and directions for future research.
@@ 2 MLB Data
We consider batter and pitcher data at the plate appearance- and pitch-level from the 2023 and 2024 MLB seasons. All regular season games were included; postseason games and spring training data were excluded.
For the stabilization analysis in Section 3, plate appearance data was obtained from FanGraphs. A plate appearance (PA) refers to a batter’s completed turn at bat, while an at-bat (AB) is a plate appearance that excludes outcomes that do not count towards batting average, i.e., walks, hit-by-pitches, and sacrifices. We focus on the following batter performance metrics (aggregated across different numbers of PAs across a season):
  • Batting average (BA): hits divided by at-bats
  • On-base percentage (OBP): the percentage of a batter’s plate appearances that result in the batter reaching base
  • Walk rate (BB%): the percentage of a batter’s plate appearances that result in a walk
  • Strikeout rate (K%): the percentage of a batter’s plate appearances that result in a strikeout
  • Hit rate (H%): the percentage of a batter’s plate appearances that result in a hit
  • Homerun rate (HR%): the percentage of a batter’s plate appearances that result in a homerun
For the changepoint analysis in Sections 5-7, we obtained pitch-level data from MLB Statcast, accessed via Baseball Savant. For pitchers, we analyze fastball velocity at the pitch level, treating four-seam fastballs and sinkers as distinct pitch types. For batters, we analyze plate discipline metrics, which evaluate a player’s ability to be selective with their swings, i.e., to swing at pitches they can hit well while taking pitches they cannot. For example, some plate discipline metrics evaluate a player’s ability to swing at pitches within the strike zone and refrain from swinging at pitches outside of the strike zone. The strike zone is defined as the area over home plate from approximately the batter’s knees to the midpoint of their torso, as measured by Statcast. In this paper, we focus on the following batter plate discipline metrics:
@@
3 Benchmarking Metric Stability Under No Change
In sports, the idea of metric “stabilization” is often discussed informally, with many practitioners deriving general thresholds from retrospective player data (Carleton 2012, Foley 2024). However, these heuristics can be misleading, as they typically assume that all observed variation reflects sampling noise, even when a player’s underlying skill may have changed. In this section, we formalize the concept of stabilization by deriving principled thresholds for when a player’s observed performance metric is likely to fall within a specified range of their true value, under the assumption that no change has occurred.
While the no change assumption can be examined using changepoint detection methods introduced later in the paper, the stabilization thresholds derived here offer a complementary diagnostic: they characterize the amount of variation one should expect due to sampling alone. This benchmark allows analysts to better interpret fluctuations, identifying when a change is statistically plausible, and also recognizing when surprising values may simply reflect limited sample size. In later sections, we characterize how frequently player performance metrics do exhibit meaningful changes and how often these changes occur in-season versus during the off season.
3.1 Metric Stabilization
Suppose a player has true underlying value $\theta$ for some metric which we estimate at time $t$ as $T_{t}$. We say that $T_{t}$ has stabilized when it becomes sufficiently close (i.e., within some threshold $S$) to $\theta$. Furthermore, we define the stabilization point as follows:
Definition 3.1 (Stabilization point)
Let a statistic $T_{t}$ be an estimate for $\theta$ at time $t$. We define the stabilization point for $T_{t}$ as the smallest value of $t$ such that the mean-squared error is bounded by $S^{2}$, i.e., $\mathbb{E}[(T_{t}-\theta)^{2}]\leq S^{2}$.
In other words, we are looking for the time point where on average we expect our estimate to be sufficiently close (i.e., within $S$) to our true parameter value.
Because player performance can change over the course of the season (see Sections 5 and 6) we want to define stabilization points for metrics without using player data when we are unsure whether or not a change occurs. Therefore, we will define stabilization points using the distributional properties of the metrics. We also discuss how to construct sequentially valid confidence intervals for metrics throughout the course of the season.
3.2 Bounded Metrics: Rate Statistics
Many player performance metrics in baseball are based on binary outcomes, e.g., on-base percentage (OBP), strikeout rate (K%), and homerun rate (HR%). These metrics can be modeled as averages of independent Bernoulli random variables. In this section, we derive stabilization benchmarks for such metrics by computing their mean squared error and constructing confidence intervals using classical concentration inequalities. We then apply this framework to several standard batting statistics and provide benchmark stabilization points that reflect typical levels of inter-player variability.
3.2.1 Stabilization Point
Consider a player skill, such as home run rate, with true underlying probability $p$ of success. Each event (e.g., plate appearance, at-bat) follows a Bernoulli distribution: $X_{i}\sim\text{Bernoulli}(p).$
The observed rate over $n$ events is given by the sample mean $\hat{p}=\sum_{i}X_{i}/n$ which has variance $Var(\hat{p})=\frac{p(1-p)}{n}$. Since $\hat{p}$ is an unbiased estimate of $p$, the mean-squared error (MSE) reduces to the variance:
$\mathbb{E}[(\frac{\sum_{i}X_{i}}{n}-p)^{2}]=Var(\frac{\sum_{i}X_{i}}{n})=\frac{p(1-p)}{n}$
To find the stabilization point, according to Definition 3.1 we find the smallest $t$ such that
$\frac{p(1-p)}{t}\leq S^{2}\implies t\geq\frac{p(1-p)}{S^{2}}.$
The number of observations required for stabilization depends not only on the precision threshold $S$, but also on the underlying skill level $p$. Stabilization occurs more quickly when $p$ is near 0 or 1, and is slowest when $p=0.5$, where the variance is maximized.
A natural way to set $S$ is as the latent standard deviation of true talent across players, i.e., the true spread of $p$ in the population. The observed variance of a metric across players, $\sigma^{2}_{\text{obs}}$, combines two sources of variation: 1) the true variance of $p$ across players, $\sigma^{2}_{\text{latent}}$, and 2) the variance due to finite-sample noise for each player’s estimate, $\sigma^{2}_{\text{samp}}$. These relate by
$\sigma^{2}_{\text{obs}}=\sigma^{2}_{\text{latent}}+\sigma^{2}_{\text{samp}}$
Therefore, we can estimate S to be one latent standard deviation:
$S=\sqrt{\sigma^{2}_{\text{obs}}-\sigma^{2}_{\text{samp}}}$
Where we define $\sigma_{\text{obs}}$ as the observed standard deviation of $\hat{p}$ across all players in the population and $\sigma_{\text{samp}}$ as the average sampling standard deviation expected due to finite $n$. Defining $n_{\text{stable}}$ as the number of observations required for the root mean squared error (RMSE) of $\hat{p}$ to fall within one latent standard deviation $\sigma_{\text{latent}}$, we have:
$n_{stable}=\frac{\hat{p}(1-\hat{p})}{\hat{\sigma}_{\text{latent}}^{2}}$
Table 1 gives the stabilization values $n_{stable}$ for various batter statistics.
Table 1: Stabilization point $n_{stable}$ for batter statistics: batting average (BA), on-base percentage (OBP), walk rate (BB%), strikeout rate (K%), hit rate (H%), and homerun rate (HR%). For each metric, the average season-ending value across players with more than 3162=4863*162=486 plate appearances is given, as well as the corresponding standard deviation values.
Metric Average $\hat{p}$ $\sigma_{obs}$ $\sigma_{samp}$ $\sigma_{latent}$ $n_{stable}$
BA .258 .030 0.019 0.016 732 AB
OBP .328 .025 0.019 0.022 442 PA
BB% 0.086 0.026 0.011 0.024 141 PA
K% 0.210 0.052 0.017 0.050 67 PA
H% 0.230 0.024 0.017 0.016 677 PA
HR% 0.034 0.013 0.007 0.011 291 PA
3.2.2 Confidence Intervals
Another way to understand statistic stabilization is through the lens of sequentially valid confidence intervals for player metrics. One straightforward approach is to construct these intervals using Hoeffding’s inequality. Since the outcome of each event (e.g., plate appearance) is modeled as $X_{i}\sim\text{Bernoulli}(p)$, we know that $0\leq X_{i}\leq 1$. By Hoeffding’s inequality, for any $v>0$, we have:
$\mathbb{P}(|\bar{X}_{t}-p|\geq t)\leq 2\exp(-2tv^{2}),$
where $\bar{X}_{t}$ is the sample mean after $t$ observations.
Setting $v=\sqrt{\frac{1}{2t}\log(2/\alpha)}$ for a desired confidence level $1-\alpha$, we obtain:
$\mathbb{P}\left(|\bar{X}_{t}-p|\geq\sqrt{\frac{1}{2t}\log(2/\alpha)}\right)\leq\alpha.$
Thus, a sequentially valid $(1-\alpha)$ confidence interval at time $t$ is given by:
$\left[\hat{p}-\sqrt{\frac{1}{2t}\log(2/\alpha)},\ \hat{p}+\sqrt{\frac{1}{2t}\log(2/\alpha)}\right].$
This framework connects naturally to our definition of stabilization: we can interpret a metric as stabilized once this confidence interval becomes narrower than a predefined tolerance. In contrast to RMSE-based thresholds, this approach provides probabilistic guarantees that remain valid at every point in time, making it well-suited for in-season monitoring of player metrics. Figure 1 plots Aaron Judge’s OBP after each game in the 2024 season, along with sequentially valid confidence intervals computed using Hoeffding’s inequality.
Figure 1: Sequentially valid confidence sequence for Aaron Judge’s OBP in the 2024 season. Each point corresponds to his OBP after a game (with the number of cumulative PAs on the x-axis). The error bars correspond to the 95% confidence interval constructed using Hoeffding’s inequality.
3.3 Extensions to other bounded metrics
While the examples above focus on binary outcomes modeled with Bernoulli distributions, many other baseball metrics are similarly bounded and can be treated using analogous methods. For instance, metrics such as average fastball velocity, launch angle, and exit velocity are continuous-valued but fall within well-defined ranges. In these cases, we can still apply concentration inequalities like Hoeffding’s by leveraging the fact that the data are bounded between known minimum and maximum values. This allows us to derive sequential confidence intervals and stabilization thresholds without making strong distributional assumptions.
For example, suppose we know that $X_{i}\in[a,b]$, then applying Hoeffding’s inequality yields:
$\mathbb{P}(|\bar{X_{t}}-\mu|\geq v)\leq 2exp(-\frac{2tv^{2}}{(b-a)^{2}})$
If we set $v=\sqrt{\frac{(b-a)^{2}}{2t}\log(2/\alpha)}$ we get that $\mathbb{P}(|\bar{X_{t}}-\mu|\geq v)\leq\alpha$. Thus, a sequentially valid confidence interval at time $t$ is $\bar{X_{t}}-\sqrt{\frac{(b-a)^{2}}{2t}\log(2/\alpha)},\bar{X_{t}}+\sqrt{\frac{(b-a)^{2}}{2t}\log(2/\alpha)}]$.
While Hoeffding-based intervals are a natural choice for bounded metrics, other valid approaches exist. For example, one can construct sequential confidence intervals using betting martingales (Waudby-Smith & Ramdas 2024), which offer nonparametric guarantees and are tightly connected to recent advances in online inference. Alternatively, one might use Gaussian intervals with corrections for multiplicity to maintain valid coverage over time.
@@ 4 Changepoint detection
Here, we are primarily concerned with the problem of identifying changes in the true mean of a given player’s performance metric. For example, detecting when a batter’s out-of-zone swing percentage decreases or when a pitcher’s fastball velocity increases.
@@ 4.1 Likelihood-Ratio Based Tests
A common method for detecting changepoints uses a likelihood-based framework (Hinkley 1970, Hinkley & Hinkley 1970). Consider a time-ordered sequence of data $y_{1:n}=(y_{1},\dots,y_{n})$. Let $p(y\mid\theta)$ denote the probability density function of $y$ given parameter $\theta$, and let $\hat{\theta}$ denote the maximum likelihood estimate for the full dataset under the null hypothesis of no changepoint.
@@
If $\lambda>c$ for some threshold $c$ (chosen to control the Type I error rate), we declare $t^{\prime}$ to be a changepoint.
Choosing an appropriate value of $c$ can be challenging in practice (Shang et al. 2017). This issue frequently arises in sports applications, as common changepoint detection algorithms often overflag changepoints, leading to high false positive (Type I error) rates (see Sections 5 and 6 for examples). In a player performance context, some false positives may be acceptable to maintain power to detect true changes, since flagged changes would typically be reviewed manually by teams. However, excessive false positives can overwhelm limited time and resources for review. In practice, generic algorithms often produce so many changepoints that manual vetting becomes infeasible.
4.2 Split sample inference
To appropriately control the type I error rate while still maintaining computational tractability, we propose combining changepoint detection with split-sample inference (Cox 1975). Split-sample inference is a statistical technique where data is divided into two parts: one part to train on (i.e., detect changepoints) and the other part to test (i.e., test whether the average difference in the metric before and after the changepoint is significant). In our setting, we split the data into even and odd indices to ensure comparable sample sizes and reduce dependence between the training and testing sets. The odd-indexed observations will be used to calculate the potential changepoint location $t^{\prime}$, and the even-indexed observations will be used to test the null hypothesis $\theta_{0}=\theta_{1}$ (or, as described in Section 6, a shifted null $\theta_{1}=\theta_{0}+\Delta$ for a pre-specified $\Delta$). The single changepoint detection algorithm, incorporating likelihood-ratio based test statistics and split sample inference, is presented in Algorithm 1.

### v2

@@ Abstract
We present a tractable framework for detecting changes in performance metrics and apply these methods to Major League Baseball (MLB) batting and pitching data from the 2023 and 2024 seasons. We propose a changepoint detection algorithm that combines a likelihood-based approach with split-sample inference to better control false positives, using either nonparametric tests or tests appropriate to the underlying data distribution. These tests incorporate a shift parameter, allowing users to specify the minimum magnitude of change to detect. We demonstrate the utility of this approach across simulation studies and several baseball applications: detecting changes in batter plate discipline metrics (e.g., chase and whiff rate), identifying velocity changes in pitcher fastballs, and validating velocity changepoints against a curated quasi-ground-truth dataset of pitchers who transitioned from relief to starting roles. Our method flags meaningful changes in 91% of these ‘ground-truth’ cases and reveals that, for some metrics, more than 60% of detected changes occur in-season. While developed for baseball, the proposed framework is broadly applicable to any setting involving monitoring of individual performance over time.
Keywords: baseball, changepoint detection, sequential monitoring, split sample inference, sports analytics
@@ 1 Introduction
Aaron Judge, an outfielder for the New York Yankees, led Major League Baseball (MLB) in on-base percentage (OBP) during the 2024 season, finishing with an impressive OBP of .458 compared to the league average of .312. Early in the season, however, his OBP after his first 24 plate appearances was just .208, substantially lower than his end-of-season OBP. Intuitively, this discrepancy highlights the well-known issue that performance metrics based on small sample sizes can vary substantially, and that this variability typically decreases as sample size increases. This example naturally raises a fundamental question relevant not just in sports analytics but in any field that tracks performance over time: how can we efficiently detect statistically meaningful shifts in an observed performance metric, or determine when apparent changes are simply due to small sample variability?
@@
Across sports, sample size in relation to metric stabilization and changepoints is a widely discussed topic with researchers often asking how large a sample is needed to be reasonably confident in the corresponding results (carleton2012small_sample, twinkietown2024sample_sizes). This concern about small sample variability is foundational in statistics, with the James-Stein estimator providing a classic example in baseball: even seemingly straightforward (batting) averages can be misleadingly variable when based on limited data (efron1977stein).
@@
In this paper, we present efficient methods for detecting changepoints in both binary and continuous performance metrics at the individual player level. Although our empirical focus is on MLB data, the underlying problem, of sequential monitoring of metrics for changes, arises across many domains, including manufacturing process control (page1954continuous, shang2017nonparametric), finance (banerjee2020change, kim2022unsupervised), and climatology (reeves2007review). While changepoint detection methods have been used in prior sports analytics research, they have primarily focused on identifying changes in aggregated statistics for the average player, team, or league across seasons (chiou2016change, whalen2024empirical). Recent work has extended changepoint detection to the individual athlete setting using Bayesian, model-based segmentation approaches for correlated performance metrics (kim2021binary).
In contrast, our focus is on computationally scalable, real-time monitoring of individual players across an entire league. The scale of current sports data, characterized by hundreds of athletes and event-level metrics, creates a multiple testing problem where traditional full-sample methods often lead to over-flagging. Rather than specifying a full generative model for player performance, which can be sensitive to misspecification and computationally intensive, we combine a likelihood-based scanning step with split-sample inference, yielding a novel framework that can be applied across a wide range of performance metrics and distributional settings. We focus on pitch-level metrics, such as fastball velocity and plate discipline rates, as these are metrics that are commonly monitored by teams and are less sensitive to contextual aggregation than outcome-based statistics (such as batting average or earned run average), while still reflecting meaningful changes in observed performance. Our approach draws on ideas from the changepoint detection literature (hinkley1970inference, sen1975tests, killick2014changepoint, matteson2014nonparametric, shang2017nonparametric, jewell2022testing) but integrates them in a novel framework specifically designed for this application.
In particular, we incorporate split-sample inference with hypothesis tests that include a shift parameter, $\Delta$. By including $\Delta$, we give users explicit control over the magnitude of changes considered meaningful. This complements and extends existing R packages for changepoint detection (lindeloev2024mcp_packages), which are primarily designed for continuous, approximately normal data and do not typically incorporate split-sample inference. By explicitly supporting both continuous and binary outcomes within a split-sample framework, our approach provides a flexible alternative for applications where non-Gaussian data and false positive control are crucial considerations (e.g., tracking changes in plate discipline metrics). We demonstrate that this framework is operationally tractable for league-wide monitoring and provide open-source R functions that implement this framework efficiently at the player level.
We validate our method via simulation studies and on a constructed quasi-ground-truth dataset of pitchers who changed roles from reliever to starter during the 2023 or 2024 MLB seasons. We illustrate the broader utility of the method by detecting changes in fastball velocity and plate discipline metrics such as whiff and chase rate.
In practice, the proposed framework is designed for routine, in-season monitoring of player performance metrics. Changepoint detection can be run at regular intervals across selected metrics to flag observable shifts in performance across large numbers of players. These flags may reflect a range of underlying phenomena, including mechanical adjustments, fatigue, injury, or other changes affecting performance. By combining a fast likelihood-based scan with split-sample confirmation, the approach supports automated detection of potentially meaningful changes for subsequent manual review, providing a rigorous, scalable tool for monitoring player performance throughout the season.
@@
The paper is organized as follows. Section 2 describes the MLB dataset. Section 3 introduces the changepoint detection algorithms. Section 4 presents simulation studies evaluating Type I error control and power. We then apply the proposed methods to batter plate discipline metrics in Section 5 and to pitcher fastball velocity in Section 6. Validation using a quasi-ground-truth dataset of pitcher role transitions is presented in Section 7. Section 8 concludes with a discussion of limitations and directions for future research.
@@ 2 MLB Data
We consider pitch-level data from the 2023 and 2024 MLB seasons obtained from MLB Statcast, accessed via Baseball Savant. Pitches from all regular season games were included; postseason and spring training data were excluded.
Pitch data are ordered chronologically within each player by pitch timestamp. For batters, sequences are constructed separately for each plate discipline metric using only the subset of pitches relevant to that metric (e.g., pitches outside the strike zone for chase rate, and pitches swung at for whiff rate). For pitchers, sequences consist of all fastballs of a given pitch type (four-seam or sinker) thrown by that pitcher during the season.
For pitchers, we analyze fastball velocity at the pitch level, treating four-seam fastballs and sinkers as distinct pitch types. For batters, we analyze plate discipline metrics, which evaluate a player’s ability to be selective with their swings, i.e., to swing at pitches they can hit well while taking pitches they cannot. For example, some plate discipline metrics evaluate a player’s ability to swing at pitches within the strike zone and refrain from swinging at pitches outside of the strike zone. The strike zone is defined as the area over home plate from approximately the batter’s knees to the midpoint of their torso, as measured by Statcast. In this paper, we focus on the following batter plate discipline metrics:
@@ 3 Changepoint detection
Here, we are primarily concerned with the problem of identifying changes in the mean of a player’s observed performance metric over time. For example, detecting when a batter’s out-of-zone swing percentage decreases or when a pitcher’s fastball velocity increases.
@@ 3.1 Likelihood-Ratio Based Tests
A common method for detecting changepoints uses a likelihood-based framework (hinkley1970inference-1, hinkley1970inference). Consider a time-ordered sequence of data $y_{1:n}=(y_{1},\dots,y_{n})$. Let $p(y\mid\theta)$ denote the probability density function of $y$ given parameter $\theta$, and let $\hat{\theta}$ denote the maximum likelihood estimate for the full dataset under the null hypothesis of no changepoint.
@@
In many classical changepoint procedures, $\lambda$ is compared to a threshold $c$ to determine whether a changepoint is present. Choosing a threshold is challenging in practice (shang2017nonparametric) and can lead to substantial overflagging in sports applications (see Sections 4-6).
In contrast, we do not threshold the likelihood-ratio statistic. Instead, we use the likelihood-ratio scan only to identify a candidate changepoint location $t^{\prime}$, and perform all formal inference using a split-sample confirmation step described in the next section.
3.2 Split-sample inference
To help control the Type I error rate while maintaining computational tractability, we combine changepoint detection with split-sample inference (cox1975note). Split-sample inference divides the data into two parts: one used to identify a candidate changepoint location, and a second, held-out portion, used to test whether a change in the metric is present.
In our setting, we split the data into even and odd indices to ensure comparable sample sizes and to reduce dependence between the detection and confirmation steps. The odd-indexed observations are used to calculate the candidate changepoint location $t^{\prime}$, while the even-indexed observations are used to test the null hypothesis $\theta_{0}=\theta_{1}$ (or, as described in Section 6, a shifted null $\theta_{1}=\theta_{0}+\Delta$ for a pre-specified $\Delta$). This design avoids the need to select a global threshold for the likelihood-ratio scan; instead, the scan is used only to identify a candidate changepoint location, which is then evaluated using held-out data.
While even-odd splitting reduces dependence between the detection and confirmation stages, it does not eliminate serial correlation or other forms of dependence potentially present in pitch-level data. The confirmation step assumes that dependence across the split is sufficiently weak that the held-out observations provide approximately independent information about the presence of a change. Thus, the confirmatory test should not be interpreted as exact under all realistic data-generating mechanisms, but rather as a practical mechanism for substantially reducing false positives relative to full-sample testing. In Section 4, we evaluate the robustness of this approach under mild serial dependence.
The full single-changepoint detection procedure, incorporating likelihood-ratio scanning and split-sample inference, is presented in Algorithm 1.